# Initial value space of the fourth order Painlevé system with $A_{5}^{(1)}\times A_{1}^{(1)}$ symmetry

### v2

@@ Abstract
The initial value spaces of the Painlevé equations are proposed by Okamoto. They are symplectic manifolds in which the Painlevé equations are described as polynomial Hamiltonian systems on all coordinates. In this article, we construct an initial value space of the fourth order Painlevé system with affine Weyl group symmetry of type $A_{5}^{(1)}\times A_{1}^{(1)}$.
Key Words: Higher order Painlevé system, Affine Weyl group, Initial value space.
@@ 1 Introduction
Sakai classified the discrete Painlevé equations from viewpoints of symmetries and IVSs in [11]. According to that, both symmetries and IVSs are labeled in terms of the affine root systems and the continuous Painlevé equations can be regarded as continuous flows commuting with some discrete Painlevé equations. On the other hand, several higher order generalizations of the Painlevé equations were proposed in [2, 3, 9, 14, 18, 22]. Hence we expect that there exists a classification theory of them from viewpoints of symmetries and IVSs.
Some IVSs of the higher order Painlevé systems have been investigated as follows.
@@
Kimura constructed the IVS of the Garnier system in $n$ variables in [4] and Suzuki constructed those of the degenerate Garnier systems in $2$ variables in [16]. In both works, the blowing up method is utilized. Tahara constructed the IVS of a fourth order Noumi-Yamada system by using Laurent series expansions of solutions in [19]. Takenawa and Carstea constructed the IVSs of some systems by resolving singularities of symmetries and gave geometric interpretations in terms of the Néron-Severi bilattices. Sasano derived those systems from the holomorphy properties.
@@
In this article, we construct an IVS of the fourth order FST system. This system is derived from a similarity reduction of the Drinfeld-Sokolov hierarchy ([2]) or the UC hierarchy ([22]) and is expressed as the polynomial Hamiltonian system
@@
$\begin{split}&\frac{dq_{i}}{dt}=\frac{\partial H}{\partial p_{i}},\quad\frac{dp_{i}}{dt}=-\frac{\partial H}{\partial q_{i}}\quad(i=1,2),\\ &H=H_{\rm VI}\left[\alpha_{3}+\alpha_{5}-\eta,\alpha_{0},\alpha_{2}+\alpha_{4},\alpha_{1}\eta;q_{1},p_{1}\right]+H_{\rm VI}\left[\alpha_{1}+\alpha_{5}-\eta,\alpha_{0}+\alpha_{2},\alpha_{4},\alpha_{3}\eta;q_{2},p_{2}\right]\\ &\qquad+\frac{(q_{1}-1)(q_{2}-t)\left\{(q_{1}p_{1}+\alpha_{1})p_{2}+(q_{2}p_{2}+\alpha_{3})p_{1}\right\}}{t(t-1)},\end{split}$ (1.1)
@@
Here the independent variable $t$ is defined in $B=\mathbb{C}\setminus\{0,1\}$. The symmetry of system (1.1) was first given in [22]. Afterward, it was shown that the group of the symmetry is isomorphic to an extended affine Weyl group of type $A_{5}^{(1)}\times A_{1}^{(1)}$ in [6]. According to that, this group is generated by the birational transformations $r_{0},\ldots,r_{5},r^{\prime}_{0},\pi_{1},\pi_{2},\rho$. Their actions on the parameters are given as follows.
@@
This article is organized as follows. In Section 2, we construct a compactification of the phase space of system (1.1) called the Hirzebruch manifold following the previous work [4, 16]. In Section 3, we determine the accessible singularities in the Hirzebruch manifold. As a result, two types of the accessible singularities $C_{j}^{(kl)}$ and $P_{j}^{(kl)}$ are obtained. In Section 4, we resolve $C_{j}^{(kl)}$ in two ways, the blowing up method and utilizing formal power series. In Section 5, we resolve $P_{j}^{(kl)}$ by using formal power series and the symmetry. In Section 6, we state the main result and consider the holomorphy property.
@@ 2 Compactification of the initial phase space
In this section, we construct a four dimensional Hirzebruch manifold $\Sigma_{\eta}$ as a compactification of the phase space $\mathbb{C}^{4}$ of system (1.1) following the previous work [4, 16].
Let $\xi$ and $\zeta^{(i)}\ (i=0,1,2)$ be homogeneous coordinates of $\mathbb{P}^{2}$ defined by
$\xi=\left[\xi_{0}:\xi_{1}:\xi_{2}\right],\quad\zeta^{(i)}=\left[\zeta_{0}^{(i)}:\zeta_{1}^{(i)}:\zeta_{2}^{(i)}\right]$
@@
[…]
We also set
$W_{ij}=\left\{\left(\xi,\zeta^{(i)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm{|}\xi_{i}\neq 0,\ \zeta_{j}^{(i)}\neq 0\right\},\quad W_{i}=W_{i0}\cup W_{i1}\cup W_{i2}\quad(i,j=0,1,2).$
Since $\det Z_{10}=-\xi_{0}^{3}\xi_{1}^{3}$ and $\det Z_{20}=-\xi_{0}^{3}\xi_{2}^{3}$, we can define manifolds $\Sigma_{\eta}^{0}$ and $\Sigma_{\eta}$ to be the quotient spaces of $W_{00}\cup W_{10}\cup W_{20}$ and $W_{0}\cup W_{1}\cup W_{2}$ respectively by relation (2.1). Then we take affine coordinates of $W_{ij}$ as follows.
@@
$\begin{split}&\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(1)},p_{2}^{(1)}\right)=\left(\frac{1}{q_{1}},\frac{q_{2}}{q_{1}},-q_{1}(q_{1}p_{1}+q_{2}p_{2}+\eta),q_{1}p_{2}\right),\\ &\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(2)},p_{2}^{(2)}\right)=\left(\frac{q_{1}}{q_{2}},\frac{1}{q_{2}},q_{2}p_{1},-q_{2}(q_{1}p_{1}+q_{2}p_{2}+\eta)\right)\end{split}$ (2.3)
and
@@
Note that $W_{ij}\simeq\mathbb{C}^{4}$ and $W_{i}\simeq\mathbb{C}^{2}\times\mathbb{P}^{2}$. Also note that $\Sigma_{\eta}\simeq T^{*}\mathbb{P}^{2}\sqcup\left(\mathbb{P}^{2}\times\mathbb{P}^{1}\right)$ if $\eta=0$ and $\Sigma_{\eta}\simeq\mathbb{P}^{2}\times\mathbb{P}^{2}$ if $\eta\neq 0$; see [4, 16] for its detail.
@@
$\frac{dq_{1}^{(1)}}{dt}=-\frac{\left(q_{2}^{(1)}\right)^{2}p_{2}^{(1)}-q_{2}^{(1)}p_{2}^{(1)}+\alpha_{3}q_{2}^{(1)}+\alpha_{1}-\eta}{t(t-1)}.$
@@
$\frac{dQ_{1}}{dt}=-\frac{(\alpha_{1}+\alpha_{3}-\eta)(\alpha_{1}+\alpha_{2}+\alpha_{3}-\eta)q_{2}^{(1)}}{t\left(q_{2}^{(1)}-1\right)}.$
@@
The system $\mathcal{H}$ has pole singularities on $\left(\Sigma_{\eta}\setminus\Sigma_{\eta}^{0}\right)\times B$.
@@
$\frac{dq_{1}}{dt}=\frac{F_{1}^{(01)}}{p_{1}^{(01)}},\quad\frac{dq_{2}}{dt}=\frac{F_{2}^{(01)}}{p_{1}^{(01)}},\quad\frac{dp_{1}^{(01)}}{dt}=G_{1}^{(01)},\quad\frac{dp_{2}^{(01)}}{dt}=\frac{G_{2}^{(01)}}{p_{1}^{(01)}},$ (2.6)
where $F_{1}^{(01)},F_{2}^{(01)},G_{1}^{(01)},G_{2}^{(01)}\in\mathbb{C}(t)\left[q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)}\right]$. Since the numerators $F_{1}^{(01)},F_{2}^{(01)},G_{2}^{(01)}$ are not divisible by $p_{1}^{(01)}$, the singularity of system (2.6) is
@@
$\left\{\left(\xi,\zeta^{(1)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm{|}\xi_{1}\neq 0,\ \zeta_{0}^{(1)}=0\right\}\times B,\quad\left\{\left(\xi,\zeta^{(2)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm{|}\xi_{2}\neq 0,\ \zeta_{0}^{(2)}=0\right\}\times B$

### v3

# Initial value space of the four dimensional Painlevé system with $(A_{5}+A_{1})^{(1)}$ symmetry

@@ Abstract
The initial value spaces of the Painlevé equations are proposed by Okamoto. They are symplectic manifolds in which the Painlevé equations are described as polynomial Hamiltonian systems on all coordinates. In this article, we construct an initial value space of the four dimensional Painlevé system with affine Weyl group symmetry of type $(A_{5}+A_{1})^{(1)}$.
Key Words: Painlevé system, Affine Weyl group, Initial value space.
@@ 1 Introduction
Sakai classified the discrete Painlevé equations from viewpoints of symmetries and IVSs in [14]. According to that, both symmetries and IVSs are labeled in terms of the affine root systems and the continuous Painlevé equations can be regarded as continuous flows commuting with some discrete Painlevé equations. On the other hand, several generalizations of the Painlevé equations were proposed in [2, 3, 12, 18, 22, 26]. Hence we expect that there exists a classification theory of them from viewpoints of symmetries and IVSs.
Some IVSs of the generalized Painlevé systems have been investigated as follows.
@@
Kimura constructed the IVS of the Garnier system in $n$ variables in [5] and Suzuki constructed those of the degenerate Garnier systems in $2$ variables in [20]. In both works, the blowing up method is utilized. Tahara constructed the IVS of a four dimensional Noumi-Yamada system by using Laurent series expansions of solutions in [23]. Takenawa and Carstea constructed the IVSs of some systems by resolving singularities of symmetries and gave geometric interpretations in terms of the Néron-Severi bilattices. Sasano derived those systems from the holomorphy properties. In this article, we construct an IVS of the four dimensional FST system.
This article is organized as follows. In Section 2, we recall the definition of initial value space and the explicit formula of the four dimensional FST system with its symmetry. In Section 3, we construct a compact manifold of the phase space of system (2.2) named $\Sigma_{\eta}$. In Section 4, the accessible singularities on $\Sigma_{\eta}$ are determined. As a result, two types of the accessible singularities $C_{j}^{(kl)}$ and $P_{j}^{(kl)}$ are obtained. In Section 5, we resolve $C_{j}^{(kl)}$ in two ways, the blowing up method and utilizing formal power series. In Section 6, we resolve $P_{j}^{(kl)}$ by using formal power series and the symmetry. In Section 7, the main result of this article is given. Some proofs of lemmas are placed in Appendix A and B.
2 Preliminary
2.1 Initial value space
We prepare some terminologies concerning the initial value space following the previous work [5, 13]. Let $\mathcal{Q}_{0}=(\mathbb{C}^{m}\times B,\pi,B)$ be a trivial fiber space with a total space $\mathbb{C}^{m}\times B$, a base space $B\subset\mathbb{C}^{n}$ and a projection $\pi$. For each $t_{0}\in B$, an inverse image $\pi^{-1}(t_{0})$ is called a fiber of $\mathcal{Q}_{0}$. We consider a Pfaff system in $\mathcal{Q}_{0}$
$dx_{i}-\sum_{j=1}^{n}X_{ij}dt_{j}=0\quad(i=1,\ldots,m),$ (2.1)
where $X_{ij}\in\mathbb{C}(t_{1},\ldots,t_{n})[x_{1},\ldots,x_{m}]$. We always assume that the base space $B$ does not contain any pole of $X_{ij}$. We also assume that any solutions to system (2.1) can be continued meromorphically to the whole $B$. Then system (2.1) defines a foliation $\mathcal{F}_{0}$ whose leaves are in one-to-one correspondence with solutions to system (2.1).
We call a leaf of $\mathcal{F}_{0}$ which has at most one intersection with any fiber of $\mathcal{Q}_{0}$ a transversal leaf. We also call a leaf which is entirely contained in a fiber of $\mathcal{Q}_{0}$ a vertical leaf. If a foliation $\mathcal{F}_{0}$ satisfies the following two conditions, then it is said to be uniform.
  1. (i)
    All leaves of $\mathcal{F}_{0}$ are transversal leaves.
  2. (ii)
    For any path $\gamma$ in $B$, if there exists a point which is an intersection of a leaf and the fiber over the starting point of $\gamma$, then the path $\gamma$ can be lifted in the leaf.
Since system (2.1) is holomorphic, all leaves of $\mathcal{F}_{0}$ are transversal leaves. However, the foliation $\mathcal{F}_{0}$ is not uniform in general because $\mathbb{C}^{m}$ is not compact. Thus we consider a fiber space $\mathcal{Q}=(E,\pi,B)$ satisfying the following three conditions.
  1. (i)
    The fiber space $\mathcal{Q}_{0}$ is a fiber subspace of $\mathcal{Q}$.
  2. (ii)
    The foliation $\mathcal{F}_{0}$ is extended to an uniform foliation $\mathcal{F}$ in the total space $E$.
  3. (iii)
    All leaves of $\mathcal{F}$ intersect with $\mathbb{C}^{m}\times B$.
Then the fiber space $\mathcal{Q}$ is said to be a P-fibration for system (2.1) and a fiber of $\mathcal{Q}$ is said to be a initial value space of system (2.1).
In order to construct a P-fibration for system (2.1), we consider a compactification of $\mathbb{C}^{m}$, which is named $\Sigma$. Then there exists a subspace of $\left(\Sigma\setminus\mathbb{C}^{m}\right)\times B$ whose point is contained in the closure of a leaf of $\mathcal{F}_{0}$. This subspace is called accessible singularity. The condition (ii) for uniform is often not satisfied at the accessible singularity. In such a case, we have to resolve the singularity with the aid of the blowing up method.
2.2 Four dimensional FST system
@@
The four dimensional FST system is derived from a similarity reduction of the Drinfeld-Sokolov hierarchy ([2]) or the UC hierarchy ([26]) and is expressed as the polynomial Hamiltonian system
@@
$\begin{split}&dq_{i}=\frac{\partial H}{\partial p_{i}}dt,\quad dp_{i}=-\frac{\partial H}{\partial q_{i}}dt\quad(i=1,2),\\ &H=H_{\rm VI}\left[\alpha_{3}+\alpha_{5}-\eta,\alpha_{0},\alpha_{2}+\alpha_{4},\alpha_{1}\eta;q_{1},p_{1}\right]+H_{\rm VI}\left[\alpha_{1}+\alpha_{5}-\eta,\alpha_{0}+\alpha_{2},\alpha_{4},\alpha_{3}\eta;q_{2},p_{2}\right]\\ &\qquad+\frac{(q_{1}-1)(q_{2}-t)\left\{(q_{1}p_{1}+\alpha_{1})p_{2}+(q_{2}p_{2}+\alpha_{3})p_{1}\right\}}{t(t-1)},\end{split}$ (2.2)
@@
Here the independent variable $t$ is defined in $B=\mathbb{C}\setminus\{0,1\}$. The symmetry of system (2.2) was first given in [26]. Afterward, it was shown that the group of the symmetry is isomorphic to an extended affine Weyl group of type $(A_{5}+A_{1})^{(1)}$ in [8]. According to that, this group is generated by the birational transformations $r_{0},\ldots,r_{5},r^{\prime}_{0},\pi_{1},\pi_{2},\rho$. Their actions on the parameters are given as follows.
@@
Remark 2.1.
It is known that a Schlesinger system enjoys the Painlevé property (see [4, 6, 10]). On the other hand, the coordinates of system (2.2) are rational in the coordinates of a certain Schlesinger system (see [15, 26]). Hence we can see that system (2.2) enjoys the Painlevé property.
@@ 3 Compactification of the initial phase space
In this section, we construct a manifold $\Sigma_{\eta}$ which is a $\mathbb{P}^{2}$ bundle over $\mathbb{P}^{2}$ as a compactification of the phase space $\mathbb{C}^{4}$ of system (2.2) following the previous work [5, 20].
Let $\xi=[\xi_{0}:\xi_{1}:\xi_{2}]$ be homogeneous coordinate of $\mathbb{P}^{2}$ and set
$U_{i}=\left\{\xi=[\xi_{0}:\xi_{1}:\xi_{2}]\in\mathbb{P}^{2}\bigm|\xi_{i}\neq 0\right\}\quad(i=0,1,2).$
Using the homogeneous coordinate $\xi$, we define affine coordinates as
$(q_{1},q_{2})=\left(\dfrac{\xi_{1}}{\xi_{0}},\dfrac{\xi_{2}}{\xi_{0}}\right),\quad\left(q_{1}^{(1)},q_{2}^{(1)}\right)=\left(\dfrac{\xi_{0}}{\xi_{1}},\dfrac{\xi_{2}}{\xi_{1}}\right),\quad\left(q_{1}^{(2)},q_{2}^{(2)}\right)=\left(\dfrac{\xi_{1}}{\xi_{2}},\dfrac{\xi_{0}}{\xi_{2}}\right).$
They imply relations
$(q_{1},q_{2})=\left(\frac{1}{q_{1}^{(1)}},\frac{q_{2}^{(1)}}{q_{1}^{(1)}}\right)=\left(\frac{q_{1}^{(2)}}{q_{2}^{(2)}},\frac{1}{q_{2}^{(2)}}\right),\quad\left(q_{1}^{(1)},q_{2}^{(1)}\right)=\left(\frac{1}{q_{1}},\frac{q_{2}}{q_{1}}\right),\quad\left(q_{1}^{(2)},q_{2}^{(2)}\right)=\left(\frac{q_{1}}{q_{2}},\frac{1}{q_{2}}\right).$
We also define affine coordinates of direct products $W_{i0}=U_{i}\times\mathbb{C}^{2}\ (i=0,1,2)$ by
$W_{00}:(q_{1},q_{2},p_{1},p_{2}),\quad W_{10}:\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(1)},p_{2}^{(1)}\right),\quad W_{20}:\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(2)},p_{2}^{(2)}\right)$
@@
$\begin{split}&\left(q_{1}^{(1)},q_{2}^{(1)},p_{1}^{(1)},p_{2}^{(1)}\right)=\left(\frac{1}{q_{1}},\frac{q_{2}}{q_{1}},-q_{1}(q_{1}p_{1}+q_{2}p_{2}+\eta),q_{1}p_{2}\right),\\ &\left(q_{1}^{(2)},q_{2}^{(2)},p_{1}^{(2)},p_{2}^{(2)}\right)=\left(\frac{q_{1}}{q_{2}},\frac{1}{q_{2}},q_{2}p_{1},-q_{2}(q_{1}p_{1}+q_{2}p_{2}+\eta)\right).\end{split}$ (3.1)
We set a manifold $\Sigma_{\eta}^{0}$ as the quotient space of $W_{00}\cup W_{10}\cup W_{20}$ by relation (3.1). Note that the manifold $\Sigma_{\eta}^{0}$ can be regarded as a $\mathbb{C}^{2}$ bundle over $\mathbb{P}^{2}$.
Let $\zeta^{(i)}=\left[\zeta_{0}^{(i)}:\zeta_{1}^{(i)}:\zeta_{2}^{(i)}\right]\ (i=0,1,2)$ be homogenous coordinates of $\mathbb{P}^{2}$ as
$p_{i}=\frac{\zeta_{i}^{(0)}}{\zeta_{0}^{(0)}},\quad p_{i}^{(1)}=\frac{\zeta_{i}^{(1)}}{\zeta_{0}^{(1)}},\quad p_{i}^{(2)}=\frac{\zeta_{i}^{(2)}}{\zeta_{0}^{(2)}}\quad(i=1,2).$
We set
$W_{ij}=U_{i}\times\left\{\zeta^{(i)}=\left[\zeta_{0}^{(i)}:\zeta_{1}^{(i)}:\zeta_{2}^{(i)}\right]\in\mathbb{P}^{2}\bigm|\zeta_{j}^{(i)}\neq 0\right\},\quad W_{i}=W_{i0}\cup W_{i1}\cup W_{i2}\quad(i,j=0,1,2).$
Note that $W_{ij}\simeq\mathbb{C}^{4}$ and $W_{i}\simeq\mathbb{C}^{2}\times\mathbb{P}^{2}$. Then relation (3.1) is extended to those
$\begin{split}&\begin{pmatrix}\zeta_{0}^{(1)}\\ \zeta_{1}^{(1)}\\ \zeta_{2}^{(1)}\end{pmatrix}=Z_{10}\begin{pmatrix}\zeta_{0}^{(0)}\\ \zeta_{1}^{(0)}\\ \zeta_{2}^{(0)}\end{pmatrix},\quad Z_{10}=\begin{pmatrix}\xi_{0}^{2}&0&0\\ -\eta\xi_{0}\xi_{1}&-\xi_{1}^{2}&-\xi_{1}\xi_{2}\\ 0&0&\xi_{0}\xi_{1}\end{pmatrix},\\ &\begin{pmatrix}\zeta_{0}^{(2)}\\ \zeta_{1}^{(2)}\\ \zeta_{2}^{(2)}\end{pmatrix}=Z_{20}\begin{pmatrix}\zeta_{0}^{(0)}\\ \zeta_{1}^{(0)}\\ \zeta_{2}^{(0)}\end{pmatrix},\quad Z_{20}=\begin{pmatrix}\xi_{0}^{2}&0&0\\ 0&\xi_{0}\xi_{2}&0\\ -\eta\xi_{0}\xi_{2}&-\xi_{1}\xi_{2}&-\xi_{2}^{2}\end{pmatrix}\end{split}$ (3.2)
on $W_{0}\cup W_{1}\cup W_{2}$. We set a manifold $\Sigma_{\eta}$ as the quotient space of $W_{0}\cup W_{1}\cup W_{2}$ by relation (3.2). Note that the manifold $\Sigma_{\eta}$ can be regarded as a $\mathbb{P}^{2}$ bundle over $\mathbb{P}^{2}$. We can take affine coordinates of $W_{ij}$ as follows.
@@
Remark 3.1.
The manifold $\Sigma_{\eta}$ was proposed by Kimura in [5] as a higher dimensional extension of the Hirzebruch surface $\Sigma_{2}$. He called $\Sigma_{\eta}$ a Hirzebruch manifold. Although we see that the manifold $\Sigma_{\eta}$ is isomorphic to $\mathbb{P}^{2}\times\mathbb{P}^{2}$ in [5, 20], it may be incorrect. Recall that the Hirzebruch surface $\Sigma_{2}$ is obtained from $\mathbb{P}^{1}\times\mathbb{P}^{1}$ through certain blowing up and blowing down procedures. The manifold $\Sigma_{\eta}$ can be obtained from $\mathbb{P}^{2}\times\mathbb{P}^{2}$ in a similar manner. We do not state its detail here.
@@
$t(t-1)dq_{1}^{(1)}+\left\{\left(q_{2}^{(1)}\right)^{2}p_{2}^{(1)}-q_{2}^{(1)}p_{2}^{(1)}+\alpha_{3}q_{2}^{(1)}+\alpha_{1}-\eta\right\}dt=0.$
@@
$t\left(q_{2}^{(1)}-1\right)dQ_{1}+(\alpha_{1}+\alpha_{3}-\eta)(\alpha_{1}+\alpha_{2}+\alpha_{3}-\eta)q_{2}^{(1)}dt=0.$
@@
The system $\mathcal{H}$ has pole singularities on $D^{0}=\left(\Sigma_{\eta}\setminus\Sigma_{\eta}^{0}\right)\times B$.
@@
$\begin{split}&t(t-1)p_{1}^{(01)}dq_{1}-F_{1}^{(01)}dt=0,\\ &t(t-1)p_{1}^{(01)}dq_{2}-F_{2}^{(01)}dt=0,\\ &t(t-1)dp_{1}^{(01)}-G_{1}^{(01)}dt=0,\\ &t(t-1)p_{1}^{(01)}dp_{2}^{(01)}-G_{2}^{(01)}dt=0,\end{split}$ (3.6)
where
$F_{1}^{(01)}=\left[(\alpha_{1}+\eta)q_{1}^{2}+\alpha_{3}q_{1}q_{2}+\{(\alpha_{0}+\alpha_{5}-\eta)t-(\alpha_{0}+\alpha_{1}+\eta)\}q_{1}-\alpha_{3}q_{2}-(\alpha_{5}-\eta)t\right]p_{1}^{(01)}$
$\qquad\qquad+(q_{1}-1)(q_{1}+q_{2})(q_{2}-t)p_{2}^{(01)}+2q_{1}(q_{1}-1)(q_{1}-t),$
$F_{2}^{(01)}=\left[\alpha_{1}q_{1}q_{2}+(\alpha_{3}+\eta)q_{2}^{2}-\alpha_{1}tq_{1}-\{(\alpha_{3}+\alpha_{4}+\eta-1)t-(\alpha_{4}+\alpha_{5}-\eta-1)\}q_{2}-(\alpha_{5}-\eta)t\right]p_{1}^{(01)}$
$\qquad\qquad+2q_{2}(q_{2}-1)(q_{2}-t)p_{2}^{(01)}+(q_{1}-1)(q_{1}+q_{2})(q_{2}-t),$
$G_{1}^{(01)}=\alpha_{1}\eta\left(p_{1}^{(01)}\right)^{2}+\{2(\alpha_{1}+\eta)q_{1}+(\alpha_{1}+\alpha_{3})q_{2}+(\alpha_{0}-\alpha_{1}+\alpha_{5}-\eta)t-(\alpha_{0}+\alpha_{1}+\eta)\}p_{1}^{(01)}$
$\qquad\qquad+(2q_{1}+q_{2}-1)(q_{2}-t)p_{2}^{(01)}+3q_{1}^{2}-2(t+1)q_{1}+t,$
$G_{2}^{(01)}=\eta\left(\alpha_{1}p_{2}^{(01)}-\alpha_{3}\right)\left(p_{1}^{(01)}\right)^{2}$
$\qquad\qquad+\bigg[\alpha_{1}(q_{2}-t)\left(p_{2}^{(01)}\right)^{2}+\{(\alpha_{1}+2\eta)q_{1}-(\alpha_{3}+2\eta)q_{2}-\alpha_{1}t+\alpha_{2}+\alpha_{3}\}p_{2}^{(01)}-\alpha_{3}(q_{1}-1)\bigg]p_{1}^{(01)}$
$\qquad\qquad+\left\{2q_{1}q_{2}-2q_{2}^{2}-2tq_{1}+(t+1)q_{2}\right\}\left(p_{2}^{(01)}\right)^{2}+\left\{2q_{1}q_{2}+2q_{2}^{2}-(t+1)q_{1}+2q_{2}\right\}p_{2}^{(01)}.$
Let $A$ be a $4\times 5$ matrix defined by
$A=\begin{pmatrix}t(t-1)p_{1}^{(01)}&0&0&0&-F_{1}^{(01)}\\ 0&t(t-1)p_{1}^{(01)}&0&0&-F_{2}^{(01)}\\ 0&0&t(t-1)&0&-G_{1}^{(01)}\\ 0&0&0&t(t-1)p_{1}^{(01)}&-G_{2}^{(01)}\end{pmatrix}$.
Then the singularity of system (3.6) is the zero set of minor determinants of $A$ of fourth order. It is given by
@@
$\left\{\left(\xi,\zeta^{(1)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm|\xi_{1}\neq 0,\ \zeta_{0}^{(1)}=0\right\}\times B$
and
$\left\{\left(\xi,\zeta^{(2)}\right)\in\mathbb{P}^{2}\times\mathbb{P}^{2}\bigm|\xi_{2}\neq 0,\ \zeta_{0}^{(2)}=0\right\}\times B$